OpenStudy (anonymous):

molar mass of P4 @Abhisar help me please

OpenStudy (abhisar):

Atomic mass of P = ?

OpenStudy (anonymous):

31

OpenStudy (abhisar):

Molar mass of P4 will be 31*4 = 124

OpenStudy (anonymous):

In my textbook it is given 132 The actual question is Find the number of mole of phosphorous in 93.0 g of phosphorous assuming the molecular formula of phosphorous is P4

OpenStudy (abhisar):

is it 0.75 moles ?

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

Then what's the problem ? Everything seems to be fine !

OpenStudy (anonymous):

mass/molar mass 93/124=:(

OpenStudy (abhisar):

Atomic mass of P = 31, so molar mass of P4 will be 4*31=124

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

Now moles in 1 gram will be 1/124. So moles in 93 grams will be 93/124 = 0.75

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

i was making the problem more difficult for myself , lol as it was under advanced problems

OpenStudy (abhisar):

O_o :D

OpenStudy (abhisar):

hv u read collisions ?

OpenStudy (anonymous):

me not yet

OpenStudy (abhisar):

k